# Binary Event-Driven Spiking Transformer

### v1

@@ 4.2.2 Impact of CIE on different architectures
Our ablation experiments on CIFAR100 demonstrate the effectiveness of the CIE method across different Bestformer architectures. The results of these ablation experiments are presented in the Figure 5(a). The Bestformer-2-384 architecture achieved 76.75% accuracy, which was significantly enhanced to 78.23% with CIE integration. Similarly, Bestformer-4-384 exhibited improved performance from 77.77% to 79.80%, while Bestformer-6-384 achieved the highest overall accuracy of 79.98% after applying CIE, compared to its baseline of 78.05%. The consistent performance improvements across all architectures suggest that CIE effectively enhances feature representation regardless of model depth. Additionally, while deeper architectures demonstrated higher baseline performance, Bestformer-4-384 achieved the most substantial improvement (+2.03%) with CIE integration, indicating an optimal balance between model capacity and enhancement effectiveness.
Figure 5: Ablation study for CIE method on CIFAR-100. (a) Impact of CIE on model accuracy. (b) A comparative analysis of information representation capability: evaluating models with and without CIE method across variable architecture depths
@@ 4.2.3 Impact of CIE on information representation
To further validate the effectiveness of CIE, we conducted an in-depth analysis of information representation capability across different encoder blocks on CIFAR-100 datasets as shown in Figure 5(b). We conducted multiple experiments to obtain the average values and corresponding error margins for all architectures to assess the overall trend. The results demonstrate that BESTformer with CIE method consistently maintains higher representation capability to original baseline (without CIE) especially in the later encoder blocks. The improved information representation ability helped the model with 4 layer encoder blocks achieve a 2.03% performance improvement. The representation gap between them remains relatively stable across different architectures, underlining the consistent superiority of our approach. These findings support the effectiveness of our CIE method, which contributes to improved performance in downstream tasks.
@@ 5 Conclusion
This work introduces a Binary Event-Driven Spiking Transformer that significantly reduces the storage and computational demands of Transformer-based Spiking Neural Networks. To address the constrained information representation capability caused by binarization, we propose the Coupled Information Enhancement (CIE) method, which combines a reversible framework and information enhancement distillation. Extensive experiments on both static and neuromorphic datasets demonstrate that BESTformer with CIE achieves superior performance compared to other binary SNNs while maintaining high efficiency. Our work provides a promising direction for developing compact yet high-performance models for resource-constrained edge devices. Future research could explore applying this approach to other neural networks and investigating its potential in real-world neuromorphic computing applications.
6 Appendix
This appendix presents the proofs of two propositions, supplementary experimental details, and the definition of the NS-ACE indicator. In Section 6.1, the proof of Proposition 1 is provided. In Section 6.2, it is demonstrated that our method is capable of preserving information pertaining to the depth of the network. Supplementary experimental settings are detailed in Section 6.3. Lastly, the calculation method of the NS-ACE indicator is elucidated in Section 6.4.
6.1 Proof of proposition 1
Proposition 1.
Proof.
Consider a deep neural network with $n$ layers. Let $\text{X}_{l}\in\mathbb{R}^{d}$ denote the random variable representing the feature map at layer $l$, where $l\in\{1,...,n\}$ and $\text{X}_{0}$ is the input. We define the layer-wise transformation as $f_{l}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, such that $\text{X}_{l}=f_{l}(\text{X}_{l-1})$. We represent the conditional information entropy of $\text{X}_{l}$ given $\text{X}_{0}$ as
[…]
Let $f_{l}^{\prime}=f_{l}\circ f_{l-1}\circ...\circ f_{1}$, then we can conveniently express the dependency between $\text{X}_{0}$ and $\text{X}_{l}$ in a formal manner, namely $\text{X}_{l}=f_{l}^{\prime}(\text{X}_{0})$. Hence, for the conditional probability distribution function $p(\boldsymbol{x}_{l}|\boldsymbol{x}_{0})$, we have
$p(\boldsymbol{x}_{l}|\boldsymbol{x}_{0})=\begin{cases}1,&\text{if }\boldsymbol{x}_{l}=f_{l}^{\prime}(\boldsymbol{x}_{0}),\\ 0,&\text{otherwise.}\end{cases}$ (17)
Additionally, $p(\boldsymbol{x}_{l},\boldsymbol{x}_{0})=0$ when $\boldsymbol{x}_{l}\neq f_{l}^{\prime}(\boldsymbol{x}_{0})$.Therefore, $\mathcal{H}(\text{X}_{l}|\text{X}_{0})$ remains to 0 since $p(\boldsymbol{x}_{l},\boldsymbol{x}_{0})\log{p(\boldsymbol{x}_{l}|\boldsymbol{x}_{0})}=0$ always holds true, which means we can simplify our mutual information to
$\mathcal{I}(\text{X}_{0};\text{X}_{l})=\mathcal{H}(\text{X}_{l})-\mathcal{H}(\text{X}_{l}|\text{X}_{0})=\mathcal{H}(\text{X}_{l}).$ (18)
According to the Data Processing Inequality Shwartz-Ziv and Tishby (2017), we obtain:
$\mathcal{I}(\text{X}_{0};\text{X}_{0})\geq\mathcal{I}(\text{X}_{0};\text{X}_{1})\geq...\geq\mathcal{I}(\text{X}_{0};\text{X}_{n}).$ (19)
Thus, we can formalize the non-increasing trend of information entropy across network layers as
$\mathcal{H}(\text{X}_{0})\geq\mathcal{H}(\text{X}_{1})\geq...\geq\mathcal{H}(\text{X}_{n}).$ (20)
6.2 Proof of proposition 2
Proposition 2.
Proof.
Let $\text{X}_{0}=(\text{X}^{0}_{0},\text{X}^{1}_{0})$ denote the input to the network, $\text{X}_{l}=(\text{X}^{0}_{l},\text{X}^{1}_{l})$ denote the output of the $l$-th layer. In our method, we define the reversible forward mapping as $\Phi_{l}(\text{X}_{l-1}^{0},\text{X}_{l-1}^{1})=(\text{X}_{l}^{0},\text{X}_{l}^{1})$ and $v$ represent the composite transformation of $n$ reversible layers, such that $v=\Phi_{n}\circ\Phi_{n-1}\circ...\circ\Phi_{1}$. Let $v^{-1}$ denote the inverse function of $v$. Applying the Data Processing Inequality Shwartz-Ziv and Tishby (2017), we obtain:
$\mathcal{I}(\text{X}_{0};\text{X}_{0})\geq\mathcal{I}(\text{X}_{0};v(\text{X}_{0}))\geq\mathcal{I}(\text{X}_{0};v^{-1}(v(\text{X}_{0})))$ (21)
Since $\text{X}_{n}=v(\text{X}_{0})$ and $\text{X}_{0}=v^{-1}(v(\text{X}_{0}))$, This chain of inequalities collapses to an equality:
$\mathcal{I}(\text{X}_{0};\text{X}_{0})=\mathcal{I}(\text{X}_{0};v(\text{X}_{0}))=\mathcal{I}(\text{X}_{0};\text{X}_{n}).$ (22)
Combining it with Equation 18, we have $\mathcal{H}(\text{X}_{0})=\mathcal{H}(\text{X}_{n}).$ In accordance with Proposition 1, this result can be extended to the intermediate layers:
$\mathcal{H}(\text{X}_{0})=\mathcal{H}(\text{X}_{1})=...=\mathcal{H}(\text{X}_{n}).$ (23)
6.3 Experiment Details
Our model implementation utilizes PyTorch Paszke et al. (2019), SpikingJelly Fang et al. (2023), and Timm Wightman (2019) frameworks. We train our model on four 24GB RTX 4090 GPUs for static datasets and one 24GB RTX 4090 GPU for neuromorphic datasets under a Linux system.
6.3.1 ImageNet Classification
ImageNet is one of the most widely used datasets for image classification tasks. It consists of 1,000 classes, including 1.3 million training images and 50,000 validation images, with an input size of 224x224. We use a patch size of $14\times 14$ and set embedding dimensions to 512. The model is trained from scratch using batch sizes of 96 and 256 for timesteps of 4 and 1, respectively. For the first 10 epochs, we employ a warmup strategy with an initial learning rate of 1e-5. Subsequently, we use the AdamW optimizer with an initial learning rate of 5e-4 and apply cosine decay over 300 training epochs. To compare our results with other quantized methods, we employ the same data augmentation techniques as Zhou et al. (2023a, b); Yao et al. (2024b), including random augmentations, mixup, and cutmix. Following common practice in binary neural networks Rastegari et al. (2016b); He et al. (2023), we leave the first input embedding layer and the last output layer unbinarized. During training, we use ViT-B/16 as the teacher model for distillation. For inference, the final prediction is obtained by averaging the outputs of the Classification Head and the Distillation Head.
6.3.2 CIFAR Classification
CIFAR datasets are classic benchmarks for image classification tasks. They provide 50,000 training images and 10,000 test images, each with a size of 32x32. We use a patch size of $4\times 4$ and embedding dimensions of 384. Our model, containing four encoder blocks, is trained from scratch using a batch size of 128. We employ the AdamW optimizer with an initial learning rate of 5e-4 and cosine decay over 400 training epochs. The first 10 epochs serve as a warm-up period, starting with an initial learning rate of 1e-5 before ramping up to the full learning rate. We use the same data augmentation techniques as Zhou et al. (2023a); Yao et al. (2024b). We utilize ViT-B/16 on CIFAR as the teacher model and average the outputs of the Classification Head and the Distillation Head for inference.
6.3.3 Neuromorphic Dataset Classification
The CIFAR10-DVS dataset is derived from CIFAR10, selecting 9,000 training images and 1,000 test images, which are then translated and captured using a DVS camera to create the neuromorphic dataset. Images in dataset are 128x128 pixels. We process these using a shallow network of two encoder blocks with embedding dimensions of 256. Experiments are conducted with time steps of 10 and 16. Training on CIFAR10-DVS lasts 106 epochs, using the AdamW optimizer with a batch size of 16. We apply the same data augmentation techniques as Zhou et al. (2023a) for CIFAR10-DVS. A full-precision Spikingformer serves as the teacher model for distillation. For inference, we average the outputs of the Classification Head and the Distillation Head.
6.4 Neuromorphic Synaptic Arithmetic Computation Effort (NS-ACE)
NS-ACE is first proposed by Shen et al. (2024). It incorporates the neuron’s firing rate ($fr_{s}$) into its calculations to more precisely capture the energy dynamics intrinsic to neuromorphic hardware systems. This model is based on the premise that energy expenditure occurs solely during spike events. To calculate NS-ACE we need define the bitbudget($\mathrm{B}\mathrm{B}$) of model as:
$\mathrm{BB}=T\cdot b_{w}\cdot b_{s},$ (24)
where $T$, $b_{w}$, $b_{s}$ represente the time step, the bit-width of weight and the bit-width of spike, respectively. Therefore, the mathematical formulation for NS-ACE is presented as follows:
$\mathrm{N}\mathrm{S}\mathrm{-}\mathrm{A}\mathrm{C}\mathrm{E}=\sum_{w\in W,s\in S}fr_{s}\cdot n_{w,s}\cdot\mathrm{B}\mathrm{B},$ (25)
where $n_{w,s}$ represents the number of multiply-accumulate (MAC) operations for a $b_{w}$-bit number and a $b_{s}$-bit number, which are directly derived from the architecture of the neural network. $W$ and $S$ refer to the sets that include all bit-widths involved during the inference phase of the neural network. $fr_{s}$ represents the firing rate of neurons.

### v2

@@ 4.2.2 Impact of CIE on different architectures
Our ablation experiments on CIFAR100 demonstrate the effectiveness of the CIE method across different Bestformer architectures. The results of these ablation experiments are presented in Figure 5(a). The Bestformer-2-384 architecture achieved 76.75% accuracy, which was significantly enhanced to 78.23% with CIE integration. Similarly, Bestformer-4-384 exhibited improved performance from 77.77% to 79.80%, while Bestformer-6-384 achieved the highest overall accuracy of 79.98% after applying CIE, compared to its baseline of 78.05%. The consistent performance improvements across all architectures suggest that CIE effectively enhances feature representation regardless of model depth. Additionally, while deeper architectures demonstrated higher baseline performance, Bestformer-4-384 achieved the most substantial improvement (+2.03%) with CIE integration, indicating an optimal balance between model capacity and enhancement effectiveness.
Figure 5: Ablation study for CIE method on CIFAR-100. (a) Impact of CIE on model accuracy across different architectures. (b) A comparative analysis of information representation capability: evaluating models with and without CIE method across variable architecture depths.
@@ 4.2.3 Impact of CIE on information representation
To further validate the effectiveness of CIE, we conducted an in-depth analysis of information representation capability across different encoder blocks on CIFAR-100 datasets as shown in Figure 5(b). We conducted multiple experiments to obtain the average values and corresponding error margins for all architectures to assess the overall trend. The results demonstrate that BESTformer with CIE method consistently maintains higher representation capability than original baseline (without CIE), especially in the later encoder blocks. The improved information representation ability helped the model with 4 layer encoder blocks achieve a 2.03% performance improvement. The representation gap between them remains relatively stable across different architectures, underlining the consistent superiority of our approach. These findings support the effectiveness of our CIE method, which contributes to improved performance in downstream tasks.
@@ 5 Conclusion
This work introduces a Binary Event-Driven Spiking Transformer that significantly reduces the storage and computational demands of Transformer-based Spiking Neural Networks. To address the constrained information representation capability caused by binarization, we propose the Coupled Information Enhancement (CIE) method, which combines a reversible framework and information enhancement distillation. Extensive experiments on both static and neuromorphic datasets demonstrate that BESTformer with CIE achieves superior performance compared to other binary SNNs while maintaining high efficiency. Our work provides a promising direction for developing compact yet high-performance models for resource-constrained edge devices.
Acknowledgments
This work is supported in part by the National Natural Science Foundation of China (No. U2333211, U20B2063 and 62220106008), in part by the Project of Sichuan Engineering Technology Research Center for Civil Aviation Flight Technology and Flight Safety (No. GY2024-27D), in part by the Open Research Fund of the State Key Laboratory of Brain-Machine Intelligence, Zhejiang University (Grant No.BMI2400020), in part by the Shenzhen Science and Technology Program (Shenzhen Key Laboratory, Grant No. ZDSYS20230626091302006), in part by the Shenzhen Science and Technology Research Fund (Fundamental Research Key Project, Grant No. JCYJ20220818103001002) and in part by the Program for Guangdong Introducing Innovative and Enterpreneurial Teams, Grant No. 2023ZT10X044.
Contribution Statement
Honglin Cao and Zijian Zhou contribute equally to this work.